# Uncertainty Quantification Metrics for Deep Regression

### v2

@@ V-C Interpretability
NLL: The lower bound of NLL is given by the ground truth distribution. But when training for real-world deployment, the true data distribution is typically unknown. The upper bound is positive infinity.
@@
AUSE: From Sec. V-B we know that ground truth distribution may not even give the optimal lower bound 0. We argue that predicted variance can only partially explain the error. There is no clear upper bound for AUSE.
@@
To sum up, there is no reference frame for the magnitudes of NLL and AUSE if ground truth distributions are unknown. I.E., NLL and AUSE can only tell us which model is better at UQ but cannot tell us how good it is. Hence, the lack of interpretability consists of a severe limitation of applying AUSE and NLL into real-world autonomous systems.
V-D Strengths and Limitations
AUSE
@@ V-D Strengths and Limitations
  • Strengths: It offers a robust quantized measure on the correlation between predictive uncertainty and errors compared with Pearson correlation and Spearman correlation. It tells how wrong a model’s prediction is likely to be, which is vital for trust-worthy autonomous system. It is more appropriate for loss prediction [34, 35], which is another UQ approach compared to predicting full distributions.
  • Limitations: The value is not bounded thus, lacks interpretability; It may fail in tasks where homoscedastic uncertainty is dominant, which is rare in real-world applications; It does not comply with distribution prediction approach since the generating distribution does not return optimal AUSE.
Calibration Error
  • Strengths: The value is statistically bounded and thus interpretable; It requires the least amount of samples to be stable, and is thus less demanding for collecting test data; The definition is flexible to extend to general confidence intervals beyond the cumulative distribution.
  • Limitations: It requires predicting a distribution; It does not capture how wrong the predictions can be.
NLL
  • Strengths: It measures the shape of the predictive distributions, including calibration and sharpness.
  • Limitations: It fails when the ground truth distribution is multimodal but the training target is unimodal[23]; It requires largest amount of test samples to be stable compared with the other two metrics.
Spearman Correlation Limitations are covered in Sec. V-A.
V-E Concluding Remarks
Essentially calibration error and NLL measure how well the predicted probability distribution corresponds to the true data generating distribution, whereas AUSE and Spearman measure how well the predicted uncertainty can tell the magnitude of errors. Which approach is deemed most useful will naturally differ between applications. For example, if an autonomous car is driving towards a wall, it needs to steer either left or right, but the exact turning angle is less important. In such a scenario the EBR model is clearly a better choice, and calibration error or NLL are clearly more appropriate metrics. On the other hand, in a scenario where an autonomous cas has to follow a road in poor visibility, then it is likely better to keep straight and report a high uncertainty, in which case the ensemble is the better model, and AUSE is the better metric.
In conclusion, we have tested four common uncertainty assessment metrics on four synthetic datasets that are designed to illustrate properties of these metrics. Calibration error and NLL exhibit similar properties, but calibration error is the more interpretable of the two. AUSE and Spearman correlation seem more appropriate for models that predict the loss directly instead of predicting the distribution. In our experiments AUSE has consistently proved to be more robust than Spearman, and therefore we discourage the use of Spearman correlation for uncertainty quantification.
@@ V-E Concluding Remarks
Acknowledgement: This work was funded by Swedish national strategic research environment ELLIIT, grant C08. Computational resources were provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS).

### v4

@@ 5.4 Interpretability
NLL: The lower bound of NLL is given by the generating distribution, which is typically unknown when training for real-world deployment. The upper bound is positive infinity.
@@
AUSE: From Sec. 5.2 we know that the generating distribution may not even give the optimal lower bound 0. We argue that predicted variance can only partially explain the error. There is no clear upper bound for AUSE.
@@
In summary, NLL and AUSE can only tell us which model is better at UQ but cannot tell us how good it is. This lack of interpretability constitutes a severe limitation of applying AUSE and NLL into real-world autonomous systems.
5.5 Concluding Remarks
We have tested four common uncertainty assessment metrics on four synthetic datasets that are designed to illustrate properties of these metrics. We also test all the metrics on a real-world regression task, Stereo disparity. Below we summarize our conclusions:
Spearman Correlation This metric is in general unsuitable. It has stability issues for small test sets, and converges to zero as the test set size grows, see Sec. 5.1.
@@ 5.5 Concluding Remarks
AUSE This metric offers a robust quantized measure on the correlation between predictive uncertainty and errors compared with Pearson correlation and Spearman correlation. It tells us how wrong a model prediction is likely to be, which is vital for trust-worthy autonomous systems. On the down side, its value is unbounded and thus it lacks interpretability, and it may fail in tasks where homoscedastic uncertainty dominates (but this is rare in real-world applications). It may be more appropriate for loss prediction (Yoo & Kweon, 2019; Cui et al., 2024), which is another UQ approach compared to predicting full distributions.
Calibration Error This metric has a clear lower bound of 0, and is highly interpretable; It requires the least amount of samples to be stable, and can be extended to general confidence intervals Kuleshov et al. (2018). Limitations are that it requires predicting a distribution, and it does not capture how wrong the predictions can be.
NLL This measures the shape of the predictive distributions, including calibration and sharpness. It fails when the generating distribution is multimodal but the training target is unimodal (Xiong et al., 2024), and it requires the largest amount of test samples to be stable among those tested here.
Essentially CE and NLL measure how well the predicted probability distribution corresponds to the true data generating distribution, but CE is the more interpretable of the two. In contrast, AUSE and Spearman measure how well the predicted uncertainty can tell the magnitude of errors. We discourage the use of Spearman correlation for uncertainty quantification, as AUSE has consistently proved to be more robust. Which metric is most useful will naturally differ between applications.
@@
Acknowledgement: This work was funded by Swedish national strategic research environment ELLIIT, grant C08. Computations for this work were enabled by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) and the Berzerlius resource provided by National Supercomputing Centre at Linköping University and the Knut and Alice Wallenberg foundation, partially funded by the Swedish Research Council through grant agreement no. 2022-06725. This project is also supported by the Wallenberg AI, Autonomous Systems, and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.